OpenStudy (anonymous):

You shoot an arrow into the air. After 0.9 seconds the arrow has gone straight upward to a height of 28.0 m above its launch point. Ignore air resistance.(a) What was the arrow's initial speed? (b) How long did it take for the arrow to first reach a height of 14 m above its launch point?

OpenStudy (aaronq):

you can solve all these problems with the kinematic equations \(\sf v_f=v_i+a\Delta t\) \(\sf x_f=x_i+v_o\Delta t +\dfrac{1}{2}a_x(\Delta t)^2\) \(\sf v_f^2=v_i^2+2a_x\Delta x\)